### v2

# Toward a Team of AI-made Scientists for Scientific Discovery from Gene Expression Data

Abstract
Machine learning has emerged as a powerful tool for scientific discovery, enabling researchers to extract meaningful insights from complex datasets. For instance, it has facilitated the identification of disease-predictive genes from gene expression data, significantly advancing healthcare. However, the traditional process for analyzing such datasets demands substantial human effort and expertise for the data selection, processing, and analysis. To address this challenge, we introduce a novel framework, a Team of AI-made Scientists (TAIS), designed to streamline the scientific discovery pipeline. TAIS comprises simulated roles, including a project manager, data engineer, and domain expert, each represented by a Large Language Model (LLM). These roles collaborate to replicate the tasks typically performed by data scientists, with a specific focus on identifying disease-predictive genes. Furthermore, we have curated a benchmark dataset to assess TAIS’s effectiveness in gene identification, demonstrating our system’s potential to significantly enhance the efficiency and scope of scientific exploration. Our findings represent a solid step towards automating scientific discovery through large language models.
1 Introduction
In the late 1990s, Netherlands Cancer Institute scientists applied machine learning and discovered 70 predictive genes for cancer spreading (Van’t Veer et al., 2002), leading to the creation of MammaPrint, a diagnostic tool for assessing cancer risk and guiding early-stage treatment (Mook et al., 2007; Brandão et al., 2019). MammaPrint sparked a billion-dollar industry and aided numerous women in cancer diagnosis and treatment. This remarkable success highlights the immense potential of machine learning in analyzing gene expression data. The availability of gene expression databases, such as the Cancer Genome Atlas (TCGA) (Tomczak et al., 2015) and the Gene Expression Omnibus (GEO) (Clough and Barrett, 2016), opens up vast opportunities for scientists to explore disease-related genes. With these resources, researchers can potentially uncover new genes important in disease development, potentially helping a wider spectrum of people suffering from various health conditions.
Furthermore, the emerging field of personalized medicine (Hamburg and Collins, 2010; Chan and Ginsburg, 2011) highlights the need for a more careful analysis. It is important to recognize that key genes linked to diseases may vary under different physical conditions. Therefore, studies should consider a diverse set of factors like age, gender, and co-occurrences of other diseases. Incorporating these conditions into research designs can help us gain a more comprehensive understanding of the underpinnings of these diseases.
This approach holds the promise of helping a broad range of patients by understanding diseases and tailoring treatments to individual needs. However, it also comes with significant challenges, such as navigating vast gene expression datasets (Hulsen et al., 2023) and addressing potential confounding factors. Additionally, researchers need to possess technical proficiency in coding, data processing, and analysis, requiring a blend of scientific knowledge and advanced analytical skills. These complexities highlight the difficulties in leveraging data analysis to benefit patients.
Leveraging Large Language Models (LLMs) (Liu et al., 2023) as agents (Zhou et al., 2023b), this paper proposes a Team of AI-made Scientists (TAIS) to automatically simulate researchers’ work. TAIS agents will execute tasks like dataset selection, preprocessing, confounder factor correction, condition prediction, and analysis to identify disease-predictive genes under various conditions.
Given the pioneering nature and complexity of our aim, We then establish a golden standard benchmark with datasets comprising 457 disease-condition pairs to assess our TAIS method’s performance. This benchmark involves manual selection, analysis, and processing of datasets, as well as writing and executing code to identify predictive genes for disease status under various conditions.
Our evaluation demonstrates that TAIS can effectively perform intricate data analysis on genetic datasets, and its performance can be further improved through the iterative process of collaboration among agents. Our case study reveals that the genes identified by our agents are corroborated by biomedical research.
In summary, the contribution of our paper is as follows.
  • We introduced the Team of AI-made Scientists (TAIS) system, an agent system simulating scientific research activities for analyzing genes predictive of disease under various conditions.
  • Beyond standard data analysis tasks (i.e. data processing, analysis), we introduced crucial steps like confounding factor correction to minimize false discoveries and two-step regression to account for missing conditions.
  • We developed a benchmark to evaluate our TAIS method, simulating human scientists’ data analysis process and documenting errors for future reference.
2 Related Work
Figure 1: The overview of the Team of the AI-made Scientists (TASI). The illustration starts from the top right corner where the user uses the system. The question goes to the project manager. The project manager further decomposes the tasks and assigns tasks to different AI-made scientists, illustrated in the yellow area. The blue area shows the details of how the statistician analyzes the data.
2.1 Pipelines to Identify Genes Predictive of Disease Status
The pipeline to analyze gene expression data with ML begins with the selection of datasets, such as public repositories like GEO (Clough and Barrett, 2016) or TCGA (Tomczak et al., 2015). The next step is data preprocessing. This involves cleaning the data, handling missing values (Abusamra, 2013), removing entries with an empty record, and eliminating genes that are not recorded (Khondoker, 2006). This step ensures the quality and reliability of the data before analysis. Then, one can perform a basic regression analysis to identify the genes that are predictive of the disease status (Ghosh and Chinnaiyan, 2005; Wu et al., 2009). Lasso (Tibshirani, 1996) is often chosen as the model due to its ability to identify a sparse set of genes, which helps pinpoint a relatively small number of genes from thousands of possible predictors. Additional steps to correct potential confounding factors (Leek et al., 2010; Bruning et al., 2016) is usually needed for a more reliable identification of the genes, especially in the scenario where the distribution of gene expressions varies across subgroups within the data due to different background distributions, rather than the disease itself (Yu et al., 2006). Recent studies extend the discoveries pipeline to incorporate more conditions in the analysis (Yang et al., 2023b), including additional covariates in the regression model, such as patient demographics, and other possible co-occurrence of diseases (Kyalwazi et al., 2023). The inclusion of these conditions allows for the identification of gene expression patterns that are not only associated with the disease status but also modulated by these conditions. This nuanced analysis supports the development of more personalized treatment strategies by identifying how different conditions affect gene-disease relationships (Rosenquist et al., 2023).
2.2 Task Solving via LLMs as Agents
LLMs have illustrated the potential to achieve human-level intelligence (Wang et al., 2023b; OpenAI, 2023; Touvron et al., 2023a, b). Building upon this capability, subsequent research explores and enhances LLMs’ ability to solve complex tasks (Wang et al., 2022a, c; Hao et al., 2023; Yao et al., 2022). For example, CoT (Wei et al., 2022) undertake tasks through goal decomposition (Zheng et al., 2023; Feng et al., 2023; Wang et al., 2022a; Ning et al., 2023) and tackles each sub-goal sequentially. Given this ability, it’s expected that the collaboration of multiple agents could further enhance the problem-solving capacities (Wang et al., 2023c; Talebirad and Nadiri, 2023; Du et al., 2023; Wang et al., 2023a), with each agent performing specific actions with distinct expertise (Yang et al., 2023a; Talebirad and Nadiri, 2023; Dong et al., 2023). For example, MetaGPT (Hong et al., 2023) promotes collaboration among various roles of AI agents (e.g. product manager, engineer, etc) by SOPs in software development. Qian et al. (2023); Dong et al. (2023) has also demonstrated role-play in software development and code generation. Other works focus on sociological phenomena (Shapiro et al., 2023; Sumers et al., 2023; Zhou et al., 2023a; Wang et al., 2023c; Li et al., 2023). Researchers have also been trying to incorporate LLMs into scientific discovery such as in chemistry (Bran et al., 2023; Guo et al., 2023), biotechnology (Madani et al., 2023) and medical (Singhal et al., 2023; Yang et al., 2023c) areas by training or finetuning LLMs on domain-specific data. In contrast to these works, We leverage available LLMs out-of-box without any training and assemble a team of AI scientists tasked with the planning, analysis, and coding to solve scientific problems in genomics.
3 Method
3.1 Overview: the Team of AI-made Scientists
Our Team of AI-made Scientists (TAIS) consists of five roles: Project Manager, Data Engineer, Domain Expert, Statistician, and Code Reviewer, each contributing distinct expertise to solve scientific problems.
Figure 2: The collaboration between Data Engineer and Code Reviewer.
Figure 2: The collaboration between Data Engineer and Code Reviewer.
Figure 3: The collaboration between Data Engineer and Domain Expert.
For a detailed description of these roles, please refer to Appendix A. The overall collaboration scheme is illustrated in Figure 1, where three AI agents (Data Engineer, Domain Expert, and Code Reviewer) are tasked with the data preprocessing procedure, and two AI agents (Statistician and Code Reviewer) with the regression analysis part.
3.2 Collaboration among AI Scientists
We highlight two critical collaborations among AI scientists in TAIS. The first, termed program-and-review, involves the interaction between the Code Reviewer and either a Statistician or Data Engineer. Throughout the coding process, code from the Data Engineer or Statistician is reviewed by the Code Reviewer agent to ensure quality and adherence to instructions, as shown in Figure 3. This process continues until the code meets the review standards or reaches the maximum number of review rounds. The Code Reviewer evaluates the code based on its executability, error-free status, and compliance with the given instructions. Successful review allows the code to proceed; otherwise, detailed feedback is provided to the coder (e.g. Data Engineer and Statistician) for revision.
The second collaboration, Consultative Coding, involves a Data Engineer working alongside a Domain Expert who offers domain-specific knowledge. For example, in a project focusing on extracting clinical information from breast cancer data, the Data Engineer managed basic data conversions but faced challenges with acronyms. Consultation with a Domain Expert introduced a method to parse complex medical terms into the expected data type, addressing issues like identifying acronyms of the Triple Negative Breast Cancer (TNBC) subtypes. This collaborative effort led to the successful completion of missing data, underscoring the importance of teamwork and expert insight in enhancing data preprocessing.
3.3 Regression Analysis on Gene Expression Data
Analyzing gene expression data to identify significant gene factors, considering confounding variables, involves a comprehensive statistical approach due to the data’s high-dimensionality and heterogeneity.
To address the challenges of variable selection in high-dimensional data, Lasso regression is employed to isolate influential genes. The detection of confounding factors is informed by analyzing the eigenvalue gaps of the covariance matrix of input features. Confounding factors, if present, necessitate either regression with confounding factor correction or regression after confounding factor adjustment, employing a linear mixed model (LMM) (Yu et al., 2006; Lippert et al., 2011; Wang et al., 2022b) or a data transformation approach respectively.
Further, to incorporate additional conditions such as age and gender into the analysis, the residualization approach is employed to account for the effect of the condition. For cases lacking direct condition data, a two-step regression strategy is adopted. This involves using common genes between datasets to estimate missing conditions, thus enabling the integration of trait and condition data for comprehensive analysis. Detailed methodologies and statistical models are discussed in Appendix B.
4 Benchmark Creation
To streamline the evaluation of our TAIS approach and promote future research with large language models in genetic data analysis, we developed the Genetic Question Exploration (GenQEX) dataset. This benchmark dataset consists of 457 carefully selected questions, complete with a comprehensive gold standard that includes genetic datasets from public sources, preprocessing and regression analysis code, and the corresponding results. Here, we outline the process of creating this benchmark.
Question Generation
A computational biology researcher identified a list of key biomedical entities related to genetics research or public health, resulting in 65 traits classified into 9 categories. These traits were paired with either another trait or demographic attributes like “age” or “gender”, generating 4556 possible pairs. These pairs were designed to pose questions in the format: “What are the significant genes related to the trait when considering the influence of the condition?” Then, we used a set of inclusion and exclusion criteria (Appendix C) and ranked the trait-condition association of pairs based on the Jaccard similarity of related genes from the NCBI Gene database, to identify 457 pairs that are of most scientific interest, which form our benchmark’s question set. Details on these pairs are available in Appendix D.
Input Dataset
Gene expression and clinical data were obtained from the Gene Expression Omnibus (GEO) (Clough and Barrett, 2016) and the Cancer Genome Atlas (TCGA) (Tomczak et al., 2015) through the Xena platform (Goldman et al., 2020). Gene symbols related to traits were sourced from the NCBI Gene database (Brown et al., 2015). For more information on these data sources, see Appendix E.
Manual Curation
A dedicated team of four researchers within our group undertook the task of curating the question list and extracting relevant input data from public sources. A subsequent phase of manual curation involved nine computer science researchers who meticulously developed the gold standard, comprising preprocessing and regression analysis code, and the outcomes. Equipped with detailed instructions and the solutions to example questions, these researchers crafted the gold standard for all listed traits over three weeks. A computational biologist provided ongoing review to ensure the rigor of the example code and instructions.
5 Experiment
5.1 Experiment Setting
To evaluate the performance of TAIS in identifying trait-related genes, we ran our system with the raw datasets as input and compared the identified genes with those from the gold standard, using metrics such as precision, recall, and Jaccard index. We started with an end-to-end evaluation on the whole data analysis process, and then we separately analyzed the preprocessing and regression abilities of our system.
5.2 Main Results
We first report the main results in Section 5.2.1. Then, to further analyze the performance of TAIS on each sub-problem, we provide ablation analysis on regression analysis in Section 5.2.2 and data preprocessing 5.2.3.
5.2.1 Performance of TAIS System
Table 1 demonstrates the performance of TAIS on our benchmark. We respectively display the performance on single-step and two-step regression problems and the overall performance. We achieve 67.74% Success Rate, 9.20% Precision and 10.40% Recall on the single-step problems. Two-step regression is relatively more difficult, which is reflected on the performance, i.e. 42.28% Success Rate, 6.49 % Precision and 6.78% Recall. Given the difficulty of the problem, the results we achieve are impressive. A qualitative analysis in Appendix F on a specific question reveals that genes pinpointed by TAIS align with findings from biomedical studies. We also experiment with other SOTA methods, e.g. MetaGPT and AutoGPT, on our benchmarks. We find that these methods are not capable of generating bug-free code on the data preprocessing. Thus, we don’t report their results here.
Table 1: Performance of TAIS System on our benchmark. We provide performance on single-step and two-step regression tasks respectively.
Success Rate (%) Precision (%) Recall (%) Jaccard (%)
Single-step 67.74 9.20 10.40 6.73
Two-step 42.28 6.49 6.78 4.18
Overall 45.73 6.86 7.28 4.53
5.2.2 Performance of Regression
To validate the ability of TAIS to perform the regression analysis to identify genes that are predictive of disease status, we replace the TCGA and GEO datasets preprocessed by our Data Engineer with the gold standard ones. Using these preprocessed data as input, we evaluated our system on the regression analysis sub-task.
Table 2: Performance of TAIS on regression analysis of our benchmark
Budget (# reviews) Success Rate (%) Precision (%) Recall (%) Jaccard (%)
Single-step 0 0.00 0.00 0.00 0.00
1 95.16 34.60 35.99 23.90
2 95.16 34.60 35.99 23.90
Two-step 0 83.54 2.08 2.06 1.19
1 5.57 0.14 0.14 0.08
2 83.54 1.81 1.83 1.09
Overall 0 72.21 1.80 1.78 1.03
1 17.72 4.82 5.00 3.31
2 85.12 6.26 6.46 4.19
Performance Table 2 demonstrates results of TAIS on regression sub-task only. As per the table, our TAIS shows a much higher Success Rate on the regression sub-task with the preprocessed data from golden standards compared with results in Table 1. Among different budgets, we achieve the best results of 85.12% Success Rate, 6.26% Precision and 6.46% Recall on overall regression problems.
Comparison with Baselines. We adopt the recent methods on LLM agents and Code Generation and conduct experiments on our benchmark. We experiment with two settings: 1. Instruct with a general task description 2. Instruct with a detailed task description, the same as ours. We experimented with MetaGPT and AutoGPT and found that despite the feedback mechanism in respective method, the generated code leads to various persistent errors in the pipeline (such as NaN during regression) and no testable output can be generated.
Different maximum number of reviews. As discussed in Section 3.2, a Code Reviewer is recruited for any task that involves code generation. During the process, multiple rounds of program-and-review are performed. One expects that with more rounds (i.e. more budgets), the generated code will be of higher quality and with fewer bugs. We perform an analysis on the budgets and demonstrate the results under different budges in Table 2. As can be seen, with more budgets, a better Success Rate and Precision-Recall can be achieved.
Table 3: Performance of TAIS on data preprocessing
Budget (# reviews) Success Rate (%) Precision (%) Recall (%) Jaccard (%)
Single-step 0 25.80 3.06 3.03 2.37
1 32.25 2.98 3.04 2.34
2 95.20 34.60 36.00 23.90
Two-step 0 49.87 9.75 9.81 7.76
1 59.49 11.27 11.37 8.91
2 83.50 1.81 1.83 1.10
Overall 0 46.60 8.84 8.89 7.02
1 55.80 10.15 10.21 8.02
2 85.10 6.30 6.50 4.20
5.2.3 Performance of Data Preprocessing
Performance Table 3 illustrates the Data Engineer’s ability on preprocessing sub-task. Among different iterations, we achieve the optimal outcomes of 95.20% Success Rate, 34.60% Precision, and 36.00% Recall on single-step regression problems and 83.50% Success Rate, 11.27% Precision and 11.37% Recall on two-step regression task.
Different maximum number of reviews We also evaluate the performance of different review iterations on the data preprocessing task. Table 3 reveals that additional review iterations of the preprocessing task generally improve the overall Success Rate, Precision, Recall, and Jaccard index. Notably, within the single-step task, a second review significantly boosts the Success Rate to 95.20%, Precision to 34.60%, Recall to 36.00%, and the Jaccard index to 23.90%.
6 Case Study
To offer a more direct understanding of the performances of our TAIS system, we detail a case study of one particular research question here. When our system is asked with “What genes are associated with Pancreatic Cancer when considering conditions related to Vitamin D Levels?” Our system identified 20+ genes with a disease (Pancreatic Cancer) prediction cross-validation accuracy of $80\%$.
The top five genes identified are SLC11A1, SOCS1, CD207, LILRB3, and SPA17. Out of these five genes, four are implicated with Pancreatic Cancer when considering the interaction with Vitamin D Levels.
SLC11A1 has been implicated in the host’s response to pathogens and may also play a role in inflammatory diseases Awomoyi (2007). Given that inflammation is a known risk factor for pancreatic cancer, and vitamin D is involved in modulating inflammatory responses Colotta et al. (2017), SLC11A1 could be a link between vitamin D levels and inflammation-related pancreatic cancer risk.
SOCS1 is a critical regulator of cytokine signaling pathways, including those involved in immune responses and inflammation Ying et al. (2019). Vitamin D is known to modulate immune function Baeke et al. (2010) and inflammation Colotta et al. (2017), suggesting that SOCS1 could be part of the pathway through which vitamin D influences pancreatic cancer risk or progression.
CD207 is a C-type lectin receptor expressed on Langerhans cells, which are involved in immune responses in the skin but might also play roles in other types of immune responses. While the direct link between CD207, vitamin D, and pancreatic cancer is less clear, the potential connection might relate to the broader immune modulation by vitamin D and how it could affect cancer immunosurveillance.
LILRB3 is involved in the regulation of immune responses, including the inhibition of various cell signaling pathways. Vitamin D has been shown to influence the immune system Baeke et al. (2010), and alterations in LILRB3 function could potentially affect how the immune system responds to cancer cells in the context of varying vitamin D levels.
SPA17 is known for its expression in reproductive tissues and certain cancers. It may play a role in cancer cell mobility and immune evasion. Given vitamin D’s effects on immune function Baeke et al. (2010), there could be a link between vitamin D levels and the immune response to pancreatic cancer cells expressing SPA17, impacting the disease’s progression or response to therapy.
7 Conclusion
We present a transformative approach to streamline the scientific discovery process through the development of a Team of AI-made Scientists (TAIS). TAIS comprises various roles, such as Project Manager and Domain expert, Each simulated by a Large Language Model (LLM). This team collaborates to execute tasks traditionally performed by data scientists, such as data preprocessing and analysis, with a focus on identifying genes predictive of disease status. To assess the efficacy of TAIS, we curated a benchmark dataset specifically for the evaluation of its performance in this domain. Our findings demonstrate a promising direction in automating the scientific discovery process, highlighting the potential of TAIS to reduce the human effort and technical expertise required in the analysis of scientific data.
Impact Statement
This paper presents work whose goal is to advance the field of scientific discover in genomics. We hope our work can provide insights for domain experts such as geneticists and help them diagnose multiple diseases in a more personalized manner. We suggest that our model should be used under human supervision to ensure a perfect result. There are many other potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Details of the roles
Below we introduce the roles of different agents in our Team of AI Scientists (TAIS).
Project Manager First of all, a Project Manager is the initiator of the TAIS with an overall view of the scientific problem (e.g. specifications and objectives) and the full knowledge of all available AI-made scientists (e.g. their capabilities). The project manager will first decompose the problem into several sub-problems based on the capabilities of available agents. For instance, the Project Manager is aware that the two datasets, i.e. TCGA and GEO, are present in the problem. Also, in the TAIS team, the Data Engineer is capable of data preprocessing and the Statistician is capable of regression analysis. Thus the Project Manager decomposes the the task of identifying genes predictive of the disease into three sub-problems, i.e. preprocessing of TCGA, preprocessing of GEO and regression analysis, as illustrated in Figure 1.
After problem decomposition, the Project Manager will then recruit AI-made scientists and assign them to each sub-problem. As illustrated in Figure 1, Data Engineer, Domain Expert and Code Reviewer are assigned to perform the two preprocessing tasks and the Statistician along with the Code Reviewer are assigned to the regression analysis task. We detail the collaboration among agents in Section 3.2. There are two types of Statisticians, capable of performing single-step and two-step regression respectively, as detailed later. The Project Manager will decide which Statistician agent is recruited for the regression analysis based on the condition variable, i.e. if the condition is Age / Gender, then recruit a single-step Statistician, otherwise a two-step Statistician.
Data Engineer A Data Engineer is designed with skills for data engineering, i.e. data analysis, code writing, and execution. In TAIS, the Data Engineer is assigned the task of preprocessing datasets, i.e. TCGA and GEO datasets. Specific context regarding the datasets is first given to the Data Engineer including the path to the raw dataset directory, the overall research question, the trait of interest, and related function tools. Then the Data Engineer will follow the instructions to perform the preprocessing process.
Moreover, the Data Engineer is capable of coding in an interactive environment. This is particularly important because each processing step is conditioned not only the previous step but also the specific data structures and the genomics information. Thus, the Data Engineer will have to execute the code at each step, check the data, and consult the Domain Expert before entering the next step. Thus, we enable the Data Engineer the ability to execute code at any step which in turn provides feedback to the coding.
To facilitate the Data Engineer in process the data in an interactive environment, we empower the agent with code execution ability with multi-step instructions and assistance from Domain Expert. As illustrated in Figure 3, at each step, Data Engineer will program following the corresponding instruction, execute the code, gather the output and asks the Domain Expert for domain information which is used for prompting the code for the next step.
Domain Expert A Domain Expert is a professional who conducts scientific investigations and experiments to understand and improve human health. Gene expression datasets are filled with biomedical terminologies, customary abbreviations, and technical descriptions of the data collection process, which are often only understandable by experts in related fields. These experts possess the knowledge and experimental techniques required to understand the samples, variables, experimental methods, and conditions from the metadata of a cohort.
Domain Experts work closely with the Data Engineer providing support on data selection and processing with their professional background. They understand the platform information and the gene measurement techniques used to determine the relevance of a cohort to the genetic question under study, help extract gene symbol information from gene annotation data, interpret or infer the patients’ clinical information from the sample characteristics portion of the dataset, which is necessary for the statistical analysis.
Statistician Statistician agents are assigned the task of performing regression analysis on the preprocessed datasets delivered by the Data Engineer. Its objective is to identify the genes that are predictive of the disease status considering different conditions. To this end, two Statistician agents are designed to perform two categories of regression, i.e. single-step and two-step regression, as detailed in Section B. Both types of Statistician agents will follow the instruction derived from Section B to perform the regression, then analyze the output and interpret and report the results.
Code Reviewer The responsibility of the Code Reviewer agent encompasses the evaluation of code quality generated by both Statisticians and Data Engineers. At every juncture of the coding workflow, the Code Reviewer agent is tasked to review the code, as illustrated in Figure 3. The process will only proceed once the code has successfully passed the review, or if the maximum number of review rounds has been reached.
Appendix B Details of Regression Analysis on Gene Expression Data
Notations We consider $\mathbf{X}\in\mathbb{R}^{n\times p}$ as the data matrix, where $n$ represents the number of samples, and $p$ the number of gene expressions. We will use $\mathbf{y}\in\mathbb{R}^{n}$ to denote the disease status. The basic data generation process is assumed to be
$\mathbf{y}=\mathbf{X}\beta+\epsilon,$ (1)
where we use $\beta$ to denote the coefficients to estimate and $\epsilon\sim N(\mathbf{0},\mathbf{I}\sigma_{\epsilon}^{2})$.
Vanilla Case In the most vanilla case, Lasso is chosen for its effectiveness in variable selection within high-dimensional datasets, isolating the most informative genes while controlling for overfitting. Lasso Tibshirani (1996) addresses this task by solving:
$\min_{\beta}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\beta\|_{2}^{2}+\lambda\|\beta\|,$ (2)
where $\beta$ is the coefficients to estimate (non-zero coefficients mean that Lasso selects these genes), and $\lambda$ is a hyperparameter.
Confounding Factor Detection The above Vanilla case is applied when there are no confounding factors in the data. However, before that, we need first check the data to inspect whether there are confounding factors. Inspired by the Marchenko-Pastur Law Johnstone (2001) and following a previous heuristic Wang et al. (2017), we inspect the “gap” of eigenvalues of the covariance matrix $\mathbf{X}\mathbf{X}^{T}$. If there is a clear gap in the eigenvalues, we consider the data are from multiple distributions, and thus, potentially with confounding factors. More formally, the confounding factor detection will return true if $e_{i}-e_{i+1}>1/n$ holds for any $i=0,1,\dots,t$, where $e_{i}$ stands for the $i$th eigenvalues if eigenvalues have been sorted descendingly, and $t$ stands for the maximum number of eigenvalues we inspect.
After the detection of confounding factors, if the agent believes there are no confounding factors in the data, it will directly follow the vanilla case above. However, if there are confounding factors, the agent will need to decide one of the following two approaches.
  • Regression with Confounding Factor Correction
  • Regression after Confounding Factor Correction
Regression with Confounding Factor Correction The first choice is to take advantage of the impressive empirical power of linear mixed model in correct confounding factors for genetic association studies Yu et al. (2006); Lippert et al. (2011); Wang et al. (2022b). With the setup of linear mixed model, we can formulate the problem with
$\mathbf{y}=\mathbf{X}\beta+\mathbf{Z}\mathbf{u}+\epsilon,$ (3)
where $\mathbf{Z}$ is an indication of the underlying population, and $\mathbf{u}\sim N(\mathbf{0},\mathbf{I}\sigma_{u}^{2})$ and $\epsilon\sim N(\mathbf{0},\mathbf{I}\sigma_{\epsilon}^{2})$, the dimension of $\mathbf{0}$ and $\mathbf{I}$ can be inferred from the context, and $\sigma$ denotes the standard deviation of the Gaussian distribution.
We use several tricks to estimate $\beta$ and the solution for a single column of $\mathbf{X}$ (denoted by $\mathbf{x}$) is
$\beta_{\mathbf{x}}=\dfrac{\mathbf{x}^{T}(\widehat{\delta}\mathbf{I}+\mathbf{X}\mathbf{X}^{T})^{-1}\mathbf{y}}{\mathbf{x}^{T}(\widehat{\delta}\mathbf{I}+\mathbf{X}\mathbf{X}^{T})^{-1}\mathbf{x}},$
because the trick will use $\mathbf{X}$ to replace the unknown $\mathbf{Z}$ Lippert et al. (2011); Wang et al. (2022b). Also, we use $\widehat{\delta}$ to denote the estimated ratio of the two $\sigma$.
Note that linear mixed model will estimate the coefficients individually for each gene, which is not necessarily ideal for our analysis. As an alternative choice, Statistician can also use the following.
Regression after Confounding Factor Correction In order to have a multivariate method, where the coefficients of genes can be considered together, the Statistician can also choose to correct the confounding factors first. The correction is mostly built upon the theory of how linear mixed model helps to correct confounding factors Wang et al. (2022b), and achieved by the rotation of $\mathbf{X}$ as following:
$\tilde{\mathbf{X}}=(\widehat{\delta}\mathbf{I}+\mathbf{X}\mathbf{X}^{T})^{-\frac{1}{2}}\mathbf{X}$
where $\widehat{\delta}$ is again the estimation ratio of the two sigmas.
With this rotation, one can perform Lasso in the vanilla case again with $\tilde{\mathbf{X}}$ and $\mathbf{y}$ for estimation of coefficients.
Incorporation of Conditions To make sure the regression analysis will incorporate different conditions such as age and gender, we will update our basic assumption of the data generation of Eq. 1 into
$\mathbf{y}=\mathbf{X}\beta+\mathbf{C}\alpha+\epsilon$ (4)
where we use $\mathbf{C}$ to denote additional conditions such as age and gender, and use $\alpha$ to denote the corresponding coefficient effect sizes.
Note that Eq. 4 is to replace Eq. 1 as an underlying data generation assumption. All the above solutions to estimate $\beta$ for Eq. 1 still apply to Eq. 4 by incorporating $\mathbf{C}$ into Eq. 2 or Eq. 3.
Two-step Regression While the above process offers a solution to incorporate the conditions when they are known. There are many cases that we do not necessarily have the conditions in the data. In this case, we will follow the following procedure, which we name two-step regression. It uses common genes between datasets that are known for their association with the condition, to estimate the conditions that our analyses need but do not know.
For example, we are interested in studying the association between disease $d$ and gene expressions given conditions $c$. However, there are no gene expression samples collected with both $d$ and $c$ appear in the same dataset. In this case, we will identify two datasets: Dataset 1 has $\mathbf{y}^{(1)}$ of disease $d$ and the corresponding gene expression $\mathbf{X}^{(1)}$, and Dataset 2 has $\mathbf{y}^{(2)}$ of the condition $c$ and the corresponding gene expression $\mathbf{X}^{(2)}$. We also need to identify a common set of genes $\mathbf{X}_{c}^{*}$ that appear in both datasets. In addition, the selection criteria for $\mathbf{X}_{c}^{*}$ also include their established biomedical relevance to the condition.
In our two-step regression, the first step is to construct the following regression model
$\mathbf{y}^{(2)}=\mathbf{X}^{(2)}_{c}\beta_{c}+\epsilon_{c}$ (5)
$\beta_{c}$ are the coefficients indicating the effects of these genes on the condition, and $\epsilon_{c}$ is the error term.
In the second step, we first predict the missing conditions for the first dataset, with
$\mathbf{C}=\mathbf{X}^{(2)}_{c}\widehat{\beta}_{c},$
where $\widehat{\beta}_{c}$ is the estimated $\beta_{c}$ from the first step.
Then, we create a $\mathbf{X}^{(2)}_{t}$ by excluding the column of $\mathbf{X}^{(2)}_{c}$ out of $\mathbf{X}^{(2)}$. Then the remaining follows by performing the analysis with $\mathbf{y}$, $\mathbf{X}^{(2)}_{t}$, and $\mathbf{C}$.
Appendix C Criteria for manual correction of trait-condition pairs
Basically, every biomedical entity in our list can be considered a trait and paired with a condition, where the condition is either another entity in the list, or a demographic attribute “age” or “gender”. However, the below rules are applied to include and exclude certain pairs to make sure that questions formed this way are scientifically valid:
  • Entities such as language abilities, Vitamin D Levels, and bone density should only serve as the condition instead of trait;
  • Entities such as obesity and hypertension, and mental disorders like anxiety disorder and bipolar disorder should be the condition to be paired with all other traits;
  • Gender-specific entities such as prostate cancer, endometriosis, and breast cancer should not be conditioned on gender, and entities from different genders should not be paired;
  • Pairs where both the trait and condition belong to the cancer category are removed. This is because questions about genetic factors behind a cancer conditional on another type of cancer are less scientifically important.
Appendix D Traits and Conditions
Table 4: Traits organized in 9 categories, and their corresponding conditions for the questions in our GenQEX benchmark.
Type Trait Conditions
1. Cancer and Oncology-Related Disorders
Liver Cancer
Endometriosis, Age, Hypertension, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Susceptibility to Infections, Obstructive sleep apnea, Gender, Anxiety disorder
Kidney Papillary Cell Carcinoma
Endometriosis, Age, Hypertension, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Gender, Susceptibility to Infections, Crohn’s Disease, COVID-19, Obstructive sleep apnea, Anxiety disorder
Kidney Chromophobe
Gender, Age, Hypertension, Obesity, Anxiety disorder
Stomach Cancer
Endometriosis, Gender, Age, Hypertension, Obesity, Anxiety disorder
Bile Duct Cancer
Gender, Age, Obesity, Anxiety disorder, Hypertension
Bladder Cancer
Endometriosis, Glucocorticoid Sensitivity, Gender, Age, Hypertension, Vitamin D Levels, Susceptibility to Infections, Crohn’s Disease, Osteoporosis, Obstructive sleep apnea, COVID-19, Obesity, Alopecia, Anxiety disorder
2. Cardiovascular Diseases
Hypertension
Age, Breast Cancer, Lung Cancer, Gender, Prostate Cancer, Obesity, Endometriosis, Obstructive sleep apnea, Pancreatic Cancer, Vitamin D Levels, Glucocorticoid Sensitivity, COVID-19, Susceptibility to Infections, Bladder Cancer, Kidney Papillary Cell Carcinoma, Osteoporosis, Liver Cancer, Head and Neck Cancer, Esophageal Cancer, Crohn’s Disease, Thyroid Cancer, Colon and Rectal Cancer, Epilepsy, Sjögren’s Syndrome, Anxiety disorder
3. Neurological and Psychiatric Disorders
Multiple Chemical Sensitivity
Anxiety disorder, Obesity, Age, Hypertension
Anxiety disorder
Obstructive sleep apnea, Gender, Hypertension, Obesity, Age
Amyotrophic Lateral Sclerosis
Age, Hypertension, Obesity, Gender, Anxiety disorder
4. Metabolic and Endocrine Disorders
Glucocorticoid Sensitivity
Bladder Cancer, Pancreatic Cancer, Endometriosis, Hypertension, Lung Cancer, Breast Cancer, Prostate Cancer, Kidney Papillary Cell Carcinoma, Thyroid Cancer, Obesity, Vitamin D Levels, Crohn’s Disease, Liver Cancer, Osteoporosis, Esophageal Cancer, COVID-19, Obstructive sleep apnea, Susceptibility to Infections, Colon and Rectal Cancer, Anxiety disorder
Osteoporosis
Bone Density, Lung Cancer, Vitamin D Levels, Hypertension, Endometriosis, Breast Cancer, Prostate Cancer, Bladder Cancer, Age, Pancreatic Cancer, Glucocorticoid Sensitivity, Gender, Obstructive sleep apnea, Obesity, Thyroid Cancer, Psoriatic Arthritis, Head and Neck Cancer, Esophageal Cancer, Colon and Rectal Cancer, Crohn’s Disease, Anxiety disorder
Polycystic Kidney Disease
Gender, Hypertension, Obesity, Anxiety disorder
5. Genetic and Developmental Disorders
Multiple Endocrine Neoplasia Type 2
Hypertension, Anxiety disorder, Obesity
Alopecia
Psoriatic Arthritis, Endometriosis, Susceptibility to Infections, Crohn’s Disease, Bladder Cancer, Obesity, Hypertension, Anxiety disorder
Intellectual Disability
Age, Obesity, Gender, Hypertension, Anxiety disorder
Craniosynostosis
Obesity, Gender, Age
Brugada Syndrome
Anxiety disorder, Hypertension, Age, Gender, Obesity
Autoinflammatory Disorders
Psoriatic Arthritis, Endometriosis, Hypertension, Obesity, Anxiety disorder
6. Gastrointestinal and Hepatic Disorders
Crohn’s Disease
Susceptibility to Infections, Pancreatic Cancer, Breast Cancer, Lung Cancer, COVID-19, Bladder Cancer, Age, Glucocorticoid Sensitivity, Prostate Cancer, Endometriosis, Psoriatic Arthritis, Obesity, Celiac Disease, Sjögren’s Syndrome, Vitamin D Levels, Hypertension, Alopecia, Gender, Obstructive sleep apnea, Kidney Papillary Cell Carcinoma, Thyroid Cancer, Head and Neck Cancer, Osteoporosis, Anxiety disorder
Celiac Disease
Crohn’s Disease, Susceptibility to Infections, Sjögren’s Syndrome, Psoriatic Arthritis, COVID-19, Endometriosis, Gender, Obesity, Hypertension, Age, Anxiety disorder
7. Respiratory and Pulmonary Disorders
Obstructive sleep apnea
Hypertension, COVID-19, Vitamin D Levels, Obesity, Endometriosis, Prostate Cancer, Bladder Cancer, Osteoporosis, Age, Breast Cancer, Lung Cancer, LDL Cholesterol Levels, Pancreatic Cancer, Susceptibility to Infections, Sjögren’s Syndrome, Thyroid Cancer, Glucocorticoid Sensitivity, Crohn’s Disease, Anxiety disorder, Liver Cancer, Kidney Papillary Cell Carcinoma
COVID-19
Susceptibility to Infections, Obstructive sleep apnea, Hypertension, Sjögren’s Syndrome, Endometriosis, Age, Pancreatic Cancer, Crohn’s Disease, Lung Cancer, Psoriatic Arthritis, Breast Cancer, Gender, Obesity, Prostate Cancer, Bladder Cancer, Vitamin D Levels, Thyroid Cancer, Head and Neck Cancer, Glucocorticoid Sensitivity, Celiac Disease, Kidney Papillary Cell Carcinoma, Anxiety disorder
8. Rheumatological and Musculoskeletal Disorders
Psoriatic Arthritis
COVID-19, Sjögren’s Syndrome, Crohn’s Disease, Alopecia, Autoinflammatory Disorders, Susceptibility to Infections, Celiac Disease, Osteoporosis, Obstructive sleep apnea, Gender, Age, Hypertension, Obesity, Anxiety disorder
Sjögren’s Syndrome
Susceptibility to Infections, COVID-19, Psoriatic Arthritis, Endometriosis, Crohn’s Disease, Celiac Disease, Obstructive sleep apnea, Pancreatic Cancer, Hypertension, Thyroid Cancer, Vitamin D Levels, Breast Cancer, Age, Obesity, Anxiety disorder
9. Miscellaneous Traits and Conditions
Endometriosis
Pancreatic Cancer, Breast Cancer, Lung Cancer, Bladder Cancer, Hypertension, Kidney Papillary Cell Carcinoma, Susceptibility to Infections, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Head and Neck Cancer, Thyroid Cancer, Liver Cancer, COVID-19, Colon and Rectal Cancer, Esophageal Cancer, Obstructive sleep apnea, Osteoporosis, Endometrioid Cancer, Crohn’s Disease, Sjögren’s Syndrome, Alopecia, Stomach Cancer, Autoinflammatory Disorders, Celiac Disease, Anxiety disorder
Susceptibility to Infections
COVID-19, Crohn’s Disease, Endometriosis, Sjögren’s Syndrome, Age, Pancreatic Cancer, Lung Cancer, Hypertension, Breast Cancer, Bladder Cancer, Gender, Prostate Cancer, Obesity, Vitamin D Levels, Thyroid Cancer, Celiac Disease, Psoriatic Arthritis, Alopecia, Obstructive sleep apnea, Kidney Papillary Cell Carcinoma, Liver Cancer, Glucocorticoid Sensitivity, Esophageal Cancer, Anxiety disorder
Kidney stones
Gender, Hypertension, Obesity, Anxiety disorder
Underweight
Obesity, Gender, Age, Anxiety disorder, Hypertension
Appendix E Details about the data sources
GEO
The Gene Expression Omnibus (GEO) (Clough and Barrett, 2016) is a public repository that stores high-throughput gene expression data, among other types. We utilized the Entrez programming utility to systematically search the GEO database for human series data pertinent to each trait in our list, focusing on datasets with a significant sample size. Both SOFT and matrix files were downloaded for each series, with heuristic file size evaluation employed to identify datasets likely containing gene expression data. For traits yielding no results from automated searches, synonym expansion via Medical Subject Headings (MeSH) terms facilitated manual data identification.
TCGA-Xena
The Cancer Genome Atlas (TCGA) (Tomczak et al., 2015), accessible through the Xena platform Goldman et al. (2020), provides a comprehensive collection of RNAseq gene expression and clinical data across many cancer types. We extracted data for 36 traits from the TCGA cohort using the UCSC Xena platform, a repository of high-quality, cancer-related gene expression and clinical data interconnected by patient IDs.
NCBI Gene
The NCBI Gene database (Brown et al., 2015) serves as a vital resource for acquiring comprehensive information on gene sequences, functions, and their associations with diseases and conditions. For each trait, we queried the database to identify a set of gene symbols known to be associated with the trait, which is used for finding disease-disease associations for question generation, and selecting common regressors for two-step regression.
Appendix F Case study
To offer a more direct understanding of the performances of our TAIS system, we detail a case study of one particular research question here. When our system is asked with “What genes are associated with Pancreatic Cancer when considering conditions related to Vitamin D Levels?” Our system identified 20+ genes with a disease (Pancreatic Cancer) prediction cross-validation accuracy of $80\%$.
The top five genes identified are SLC11A1, SOCS1, CD207, LILRB3, and SPA17. Out of these five genes, four are implicated with Pancreatic Cancer when considering the interaction with Vitamin D Levels.
SLC11A1 has been implicated in the host’s response to pathogens and may also play a role in inflammatory diseases Awomoyi (2007). Given that inflammation is a known risk factor for pancreatic cancer, and vitamin D is involved in modulating inflammatory responses Colotta et al. (2017), SLC11A1 could be a link between vitamin D levels and inflammation-related pancreatic cancer risk.
SOCS1 is a critical regulator of cytokine signaling pathways, including those involved in immune responses and inflammation Ying et al. (2019). Vitamin D is known to modulate immune function Baeke et al. (2010) and inflammation Colotta et al. (2017), suggesting that SOCS1 could be part of the pathway through which vitamin D influences pancreatic cancer risk or progression.
CD207 is a C-type lectin receptor expressed on Langerhans cells, which are involved in immune responses in the skin but might also play roles in other types of immune responses. While the direct link between CD207, vitamin D, and pancreatic cancer is less clear, the potential connection might relate to the broader immune modulation by vitamin D and how it could affect cancer immunosurveillance.
LILRB3 is involved in the regulation of immune responses, including the inhibition of various cell signaling pathways. Vitamin D has been shown to influence the immune system Baeke et al. (2010), and alterations in LILRB3 function could potentially affect how the immune system responds to cancer cells in the context of varying vitamin D levels.
SPA17 is known for its expression in reproductive tissues and certain cancers. It may play a role in cancer cell mobility and immune evasion. Given vitamin D’s effects on immune function Baeke et al. (2010), there could be a link between vitamin D levels and the immune response to pancreatic cancer cells expressing SPA17, impacting the disease’s progression or response to therapy.